# Who Made This? Fake Detection and Source Attribution with Diffusion Features

### v1

@@ 4.2 Fake Image Detection
Our first goal is to design a fake image detection pipeline capable of generalizing to images from unseen generators. To this aim, we evaluate two approaches: a simple neural network (MLP) and a distance-based classifier ($k$-NN). In both cases, we follow the cross-generator evaluation protocol described in Section 4.1, using the selected U-Net layer.
@@
All the models have two output neurons with a softmax activation function and are trained using binary cross-entropy loss with AdamW optimizer and a learning rate of $3\times 10^{-4}$. The validation set accuracy is used to early stop the training process with a patience of 15 epochs.
@@
The best MLP and $k$-NN models are selected following the cross-generator protocol described in Section 4.1, based on validation accuracy, and then tested on the GenImage test set.
@@ 4.3 Source Model Attribution
Our second objective is source generator attribution, a multi-class classification task to identify the source model of an image. As for fake image detection, we compare the MLP and $k$-NN approaches, adapting the models and the evaluation protocol for multi-class classification. More specifically, we employ a single classifier with nine output classes (eight generators and real images).
@@
We evaluate the same MLP configurations used for fake-image detection (see Section 4.2), but replace the output layer with a 9-way softmax classifier and train with cross-entropy loss. All MLPs are trained on synthetic images from the training subsets of all generators, while real images are sampled from the training subset of the Midjourney dataset. All models are trained using the AdamW optimizer with a learning rate of $3\times 10^{-4}$, and the validation accuracy is used to early stop the training process using a patience of 15 epochs.
@@
[…]
The most effective approach is then evaluated on the GenImage test set for the attribution task.
@@ 5 Results
This section presents our experimental results, starting with the analysis of features extracted from the diffusion U-Net (Section 5.1), and then covering fake image detection (Section 5.2) and source model attribution (Section 5.3).
@@ 5.1 U-Net Layer Selection by Linear Probing
Following the experimental setup in Section 4.1, we test image prototypes from all the layers of the diffusion U-Net using a linear probing approach to determine (i) if the latent space representation can be used to distinguish between real and synthetic data, and (ii) which layer produces the most informative representation for this task. Figure 2 plots the average cross-generator validation accuracy for features extracted from different levels of the diffusion U-Net.
@@
The results of our feature-probing analysis show that the latent features extracted from the diffusion U-Net can be highly discriminative for distinguishing real from synthetic images. In particular, we identify the first layer of the decoder at $16\times 16$ resolution (Decoder 16_0) as the source of the most informative features, yielding a peak average cross-generator validation accuracy of $78.50\%$. Given this result, we adopt these features in the remainder of our study.
@@ 5.2 Fake Image Detection
Following the experimental setup described in Section 4.2, we evaluate two methods: one based on neural networks (i.e., MLP) and the other on a $k$-NN approach.
@@
We conduct preliminary experiments using the three MLPs described in Section 4.2, MLP-320, which achieves an average cross-generator accuracy of $79.17\%$, outperforms MLP-640 ($78.23\%$) and MLP640-320 ($78.94\%$). A detailed breakdown of the results for MLP-320 on cross-generator fake image detection is presented in Table 1.
Trained on Tested on
@@
The results reveal a clear pattern: while the MLP effectively learns generator-specific artifacts (as evidenced by the high diagonal accuracy), it struggles to generalize to unseen generators. This behavior suggests the presence of distinct generator families that share underlying characteristics within the latent space of the diffusion U-Net. For instance, the MLP trained on Midjourney data performs well on images from SDV1.4, SDV1.5, and Glide, indicating a latent similarity among their generated outputs.
@@
While the features extracted from the diffusion U-Net are discriminative enough to separate real and synthetic data, the MLP classifier tends to learn specific artifacts of a given generator, which limits its ability to generalize. To address this limitation, we propose a $k$-NN approach. As a training-free method that relies on distances between feature prototypes, we hypothesize that $k$-NN can achieve a more robust generalization than the MLP.
First, we identify the optimal $k$-NN configuration following the extensive parameter search described in Section 4.2. The model using the correlation distance, with $k=101$ and a support size of $2000$, achieved the best performance on the validation set (see Table 2). It yielded an average accuracy of $85.3\%$ across the eight generators, a significant improvement over the $79.17\%$ obtained by the best MLP approach. This optimized version of the $k$-NN model is finally tested on the GenImage test set (Table 3).
Support set from Tested on
Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
Midj. 88.1 87.0 86.9 84.7 88.3 88.3 87.3 86.3 87.1
SDV1.4 88.3 89.9 89.6 83.2 86.7 62.6 83.8 88.3 84.0
SDV1.5 89.4 89.7 90.4 84.4 89.8 87.1 87.2 88.7 88.3
ADM 83.0 82.5 83.5 82.4 83.6 83.7 84.5 82.6 83.2
Glide 85.7 83.9 84.1 85.4 87.4 87.4 87.8 83.8 85.7
Wukong 81.2 75.5 76.7 84.5 89.5 88.6 87.8 77.8 82.7
VQDM 83.0 83.4 84.1 83.6 84.5 84.7 85.5 83.0 84.0
BigGAN 88.1 87.8 88.5 84.2 88.3 88.7 88.7 87.7 87.7
Avg. 85.8 85.0 85.5 84.0 87.3 83.9 86.5 84.8 85.3
Table 2: Cross-generator evaluation of $k$-NN on the GenImage validation set for fake image detection. We report the accuracy of the selected $k$-NN configuration. For each generator, we use the support set from one model and test it on the validation images of all the generators.
Support set from Tested on
Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
Midj. 91.4 89.1 89.3 88.0 91.5 91.6 89.9 88.6 89.9
SDV1.4 89.9 90.9 90.7 85.4 88.2 64.1 84.7 89.4 85.4
SDV1.5 91.7 91.7 91.4 86.7 91.9 89.9 89.4 90.9 90.4
ADM 86.2 85.5 85.6 86.7 86.9 86.5 86.5 85.1 86.1
Glide 89.0 86.7 86.8 89.6 91.1 90.6 90.0 87.2 88.9
Wukong 85.3 78.7 79.5 88.2 93.0 92.7 91.2 81.6 86.3
VQDM 86.7 86.6 86.5 87.9 88.6 88.1 87.9 85.8 87.3
BigGAN 90.8 90.3 89.9 87.4 90.9 91.6 90.9 89.9 90.2
Avg. 88.9 87.4 87.5 87.5 90.3 86.9 88.8 87.3 88.1
Table 3: Cross-generator evaluation of $k$-NN on the GenImage test set for fake image detection. We report the accuracy of the selected $k$-NN configuration. For each generator, we use the support set from one model and test it on the images of all the generators.
@@
CNNSpot [46] 58.2 70.3 70.2 57.0 57.1 67.7 56.7 56.6 61.7
Spec [56] 56.7 72.4 72.3 57.9 65.4 70.3 61.7 64.3 65.1
F3-Net [33] 55.1 73.1 73.1 66.5 57.8 72.3 62.1 56.5 64.6
GramNet [26] 58.1 72.8 72.7 58.7 65.3 71.3 57.8 61.2 64.7
DIRE [49] 65.0 73.7 73.7 61.9 69.1 74.3 63.4 56.7 67.2
LaRE2 [29] 66.4 87.3 87.1 66.7 81.3 85.5 84.4 74.0 79.1
LATTE [43] 71.3 79.3 81.8 82.2 92.8 82.0 82.9 87.8 82.5
FRIDA (Ours) 88.9 87.5 87.5 87.5 90.3 86.9 88.9 87.4 88.1
Table 4: Fake image detection comparison with state-of-the-art approaches on GenImage test set. The reported metric is the average percentage accuracy calculated from eight distinct models, each trained on the data from a different generator, and tested on the specified test set. † indicates that the results are reproduced by [29]. Best in bold, second best underlined.
As detailed in Table 4, our proposed approach demonstrates a significant leap in performance over current state-of-the-art methods on the GenImage test set. The experimental setup is designed to rigorously test generalization: eight models are trained on limited data (2000 samples) from a single generator each, but evaluated on images from all generators. In this challenging scenario, our method is the one that better generalizes across these unseen data distributions, establishing a new state-of-the-art performance by a margin of nearly six percentage points.
Beyond its accuracy, the primary advantage of our approach lies in its remarkable data and training efficiency. The strong results are achieved without a conventional training process; we use a pre–trained SDM and the $k$-NN framework merely requires storing the support set. The proposed approach maintains high performance even with a drastically smaller support set. When the sample size is reduced by $90\%$, the average cross-generator accuracy on the test set drops by only $1.6\%$ (from $88.1\%$ to $86.5\%$), a result that still outperforms the previous state-of-the-art (Table 4). This characteristic is crucial for practical deployment. In a field where new generative models are constantly emerging, our method provides a scalable and effective solution that does not require cost-prohibitive data generation and retraining cycles. Using an NVIDIA RTX 4090, SDM feature extraction required an average of 2.1 seconds, while $k$-NN classification took 0.0003 seconds. In contrast, on a CPU (Intel Core i7-9800X @ 3.80 GHz), feature extraction averaged 12.9 seconds and classification 0.002 seconds.
@@ 5.3 Source Model Attribution
We compare MLP and $k$-NN classifiers on their ability to identify an image’s source generator, following the experimental procedure described in Section 4.3.
@@
A portion of the results from the hyperparameter selection on the validation set obtained using the correlation distance is presented in Table 5.
$k$ Support sizes 18 90 180 900 1800 3600 9000 1 25.2% 33.5% 36.4% 44.6% 48.1% 51.6% 55.8% 9 28.7% 31.3% 44.09% 48.7% 52.7% 57.7% 19 27.2% 42.5% 48.3% 52.1% 57.1% 35 39.5% 46.5% 50.9% 56.1% 101 39.7% 46.3% 52.0%
Table 5: Effect of $k$ and support set size on source attribution accuracy. Validation accuracies for the nine-class attribution task (eight generators plus real images) computed across different values of $k$ and support set sizes.
These results proved that the $k$-NN classifier is inadequate for effective source image attribution. Despite optimization (using a support set of 9,000 and k=9), its peak accuracy reached a mere $57.7\%$, a performance level far too low for practical application.
@@
We train the three MLP classifiers as described in Section 4.3; in Table 6 we report the average results of the MLPs on the validation set across ten training runs.
@@
Table 6: Source attribution performance of MLP models. Average accuracy and standard deviation over ten runs on the nine-class attribution task (eight generators plus real images) on the validation set.
@@
While the average accuracies of the three models on the source attribution task are comparable, the MLP-640, with $84.87\%$ of average accuracy, slightly outperforms the others and allows for an increase in performance of about 27 percentage points over the $k$-NN classifier. The MLP-640, when tested on the GenImage test set, achieves an accuracy of $84.36\%$. This result is consistent with the validation set performance, indicating a high degree of generalization.
Figure 3: Confusion Matrix for source image attribution. The MLP-640 is evaluated on the GenImage test set.
As shown in Figure 3, the classifier generally recognizes source generators with high accuracy; however, it struggles to distinguish models that share the same architecture. In particular, it frequently confuses SDM v1.4 with SDM v1.5, and, in some cases, the Wukong model — which is based on SDM — is also mistaken for them.
@@
Notably, the huge performance increase of the MLP over the $k$-NN classifier on the validation set confirms that the features extracted by the diffusion U-Net contain generator-specific patterns. The MLP is capable of learning these distinguishing characteristics, whereas the $k$-NN approach, which relies solely on feature distances, cannot recognize these subtle patterns. Although $k$-NN proves effective for the broader task of distinguishing real from synthetic content, it is insufficient for the more nuanced task of source attribution. This suggests that while the feature distance between real and synthetic images is large, the distances between features from different generators are smaller. The key differentiators are not the absolute distances but rather specific, patterns within the features that the MLP can successfully identify.
To investigate this hypothesis, we employ the SHAP algorithm to interpret the decisions of the MLP-640 classifier. We utilize the Gradient Explainer with a background dataset of 200 samples and 1000 test samples. For each class, we identify the top 10 most informative features. In Figure 4 we report the percentage of the top 10 features shared between each pair of generators.
Figure 4: Shared informative features across generators. We use SHAP with the Gradient Explainer to interpret the decisions of the MLP-640 classifier, employing 200 background and 1000 test samples. The plot reports, for each pair of generators, the percentage of overlap among their top 10 most informative features.
[Real]Refer to caption [ADM]Refer to caption [BigGAN]Refer to caption
[Glide]Refer to caption [Midjourney]Refer to caption [SDV1.4]Refer to caption
[SDV1.5]Refer to caption [VQDM]Refer to caption [Wukong]Refer to caption
Figure 5: Impact of the top-10 features on model decisions by SHAP analysis. Visualization of the ten most influential features identified by the SHAP analysis for each of the eight diffusion generators and the real image class. The figure highlights how different feature subsets contribute to the network’s decision process across generators, revealing both shared and model-specific attribution patterns.
@@
The two SDMs exhibit a $50\%$ overlap in their top 10 most important features. Furthermore, these shared features influence the model’s output identically (see Figure 5 and Figure 5), indicating that the images generated by these two models lead to a latent representation with common characteristics. As a result, the images generated by the two models are nearly indistinguishable to MLP. In contrast, the classifier successfully discriminates real and BigGAN images from the SDMs by exploiting divergent feature behavior. Although these models share $30\%$ of their top 10 features with the SDMs, the impact of these common features is markedly different (Figure 5 and Figure 5). Finally, the Wukong model presents an interesting paradox. Despite occasional misclassifications with the SDM-generated images, the SHAP analysis reveals zero feature overlap within the top 10 most influential features (Figure 5). This indicates that the classifier’s errors are not driven by the same high-impact features that define the SDM class. Instead, the confusion likely arises because a different set of features in Wukong images produces a combined effect that coincidentally mimics the characteristics of an SDM, leading the classifier to an incorrect conclusion.

### v2

@@ 4.2 Fake Image Detection
Our first goal is to design a fake image detection pipeline capable of generalizing to images from unseen generators. To this aim, we evaluate two approaches: a simple neural network (MLP) and a distance-based classifier ($k$-NN). In both cases, we extract the features from the selected U-Net layer and follow the cross-generator evaluation protocol described in Section˜3.2 to select the best model configurations based on the validation accuracy. The best MLP and $k$-NN models are then tested on the GenImage test set.
@@
All the models have two output neurons with a softmax activation function and are trained using binary cross-entropy loss with AdamW optimizer and a learning rate of $3\times 10^{-4}$. The validation accuracy is used to early stop the training process with a patience of 15 epochs.
@@ 4.3 Source Model Attribution
Our second objective is image attribution, a multi-class classification task to identify the source model of an image. As for fake image detection, we compare the MLP and $k$-NN approaches, adapting the models and the evaluation protocol for multi-class classification. More specifically, we employ a single classifier with nine output classes (eight generators + real images).
@@
We evaluate the same MLP configurations used for fake-image detection (see Section˜4.2), but we replace the output layer with a 9-way softmax classifier and train with cross-entropy loss. All MLPs are trained on synthetic images from the training subsets of all generators, while the real images are sampled from the training subset to maintain the class balance. All models are trained using the AdamW optimizer with a learning rate of $3\times 10^{-4}$, and the validation accuracy is used to early stop the training process using a patience of 15 epochs.
@@
We perform the hyperparameter selection for $k$-NN using the same grid employed in Section˜4.2, changing the sizes of the support set $S$ (ranging from 18 to 9,000) [^4] Sizes(S)=Sizes(S)={18, 45, 90, 180, 360, 450, 900, 1350, 1800, 2700, 3600, 4500, 9000} to comprise an equal number of images for all nine classes.
4.4 Extended Experimental Analysis
Benchmarking Against ViT Backbones.
We assess the advantage of using SDM features in our method by conducting a comparative analysis against established ViT-based feature extractors, specifically CLIP and DINO. For both, we extract the features from their last layer (widely used for their robust performance across various downstream tasks) and we evaluate the cross-generator performance under the same setup used for the SDM.
Robustness to Image Corruptions.
We evaluate the robustness of FRIDA by introducing various synthetic distortions to the test set. Specifically, we assess the cross-generator performance across multiple degradation levels, including JPEG compression (10, 30, 50, 70, 90), gaussian noise ($\sigma$=0.05, 0.1, 0.15, 0.2, 0.25), and gaussian blur ($\sigma$=1, 2, 3, 4, 5).
Out-of-Distribution Robustness.
We further evaluate the generalization capabilities of our approach on OOD datasets. The support set for $k$-NN is a random subset of the GenImage dataset encompassing all available generators and their corresponding real images. During inference, we evaluate the model against a series of unseen OOD generators. To isolate the domain shift introduced by the synthetic data, the real images in the test set remain constant across all evaluations, comprising 500 samples balanced equally between Pascal VOC ($N=250$) and LAION-400M ($N=250$). For each test iteration, these are paired with 500 synthetic images generated by a single OOD model, thereby maintaining a strictly balanced binary classification setup.
@@ 5 Results
This section presents our experimental results, starting with the U-Net layer selection for feature extraction (Section˜5.1), and then covering fake image detection (Section˜5.2) and source model attribution (Section˜5.3). Finally, in Section˜5.4, we present the results of the extended analysis using different backbones, OOD datasets and various image perturbations.
@@ 5.1 U-Net Layer Selection by Linear Probing
Following the experimental setup in Section˜4.1, we test image prototypes from all the layers of the diffusion U-Net using a linear probing approach to determine (i) if the latent space representation can be used to distinguish between real and synthetic data, and (ii) which layer produces the most informative representation for this task. Figure˜2 plots the average cross-generator validation accuracy for features extracted from different levels of the diffusion U-Net.
@@
The results of our feature probing analysis show that the latent features extracted from the diffusion U-Net can be highly discriminative for distinguishing real from synthetic images. In particular, we identify the first layer of the decoder at $16\times 16$ resolution (Decoder 16_0) as the source of the most informative features, yielding a peak average cross-generator validation accuracy of $78.50\%$. Given this result, we adopt these features in the remainder of our study.
@@ 5.2 Fake Image Detection
Following the experimental setup described in Section˜4.2, we evaluate two methods: one based on neural networks (i.e., MLP) and the other on a $k$-NN approach.
@@
We conduct preliminary experiments using the three MLPs described in Section˜4.2. MLP-320, which achieves an average cross-generator accuracy of $79.1\%$, outperforms MLP-640 ($78.2\%$) and MLP-640-320 ($78.9\%$). A detailed breakdown of the results for MLP-320 on cross-generator fake image detection is presented in Table˜1.
Trained on Accuracy (%) when tested on
@@
The results reveal a clear pattern: while the MLP effectively learns generator-specific artifacts (as evidenced by the high diagonal accuracy), it struggles to generalize to unseen generators. We can also observe the presence of distinct generator families that share underlying characteristics within the latent space of the diffusion U-Net. For example, an MLP trained on SDM latents of Midjourney images transfers well to SDv1.4, SDv1.5, and Glide, suggesting a strong similarity in their latent feature representations.
@@
Although U-Net features effectively distinguish real from synthetic data, the MLP classifier tends to learn generator-specific artifacts, hindering generalization. To address this limitation, we propose a training-free $k$-NN approach that operates on distances between feature prototypes. By relying on the intrinsic geometry of the feature space rather than learned parameters, $k$-NN leverages the structure already encoded in the representations, leading to more effective cross-generator generalization. To this aim, guided by the hyperparameter search described in Section˜4.2, we select the $k$-NN model using correlation distance with $k=45$ and a size of $S_{G}=1000$. This configuration achieves the best validation accuracy, $85.3\%$ across eight generators as shown in Table˜2(a), and surpasses the best MLP ($79.1\%$). The same $k$-NN configuration is then evaluated on the GenImage test set (Table˜2(b)).
(a) Results on GenImage Validation set
$S_{G}$ Accuracy (%) when tested on Midj. SDv1.4 SDv1.5 ADM Glide Wukong VQDM BigGAN Avg. Midj. 88.0 85.8 86.3 83.8 87.9 87.6 86.1 85.1 86.3 SDv1.4 88.1 90.0 90.7 83.5 87.7 65.9 84.7 88.1 84.8 SDv1.5 89.0 89.2 89.7 83.0 88.6 85.5 86.6 88.2 87.5 ADM 83.5 82.4 83.6 82.6 84.0 84.2 84.5 82.6 83.4 Glide 85.9 84.4 84.5 85.7 87.8 87.8 88.5 84.3 86.1 Wukong 82.1 74.8 75.7 84.5 89.7 89.0 88.5 76.8 82.6 VQDM 82.7 83.2 83.8 83.9 85.1 85.2 85.9 83.6 84.2 BigGAN 86.9 87.3 88.3 83.4 88.3 88.6 88.4 87.3 87.3 Avg. 85.8 84.6 85.3 83.8 87.4 84.2 86.6 84.5 85.3
(b) Results on GenImage Test set
$S_{G}$ Accuracy (%) when tested on Midj. SDv1.4 SDv1.5 ADM Glide Wukong VQDM BigGAN Avg. Midj. 91.1 88.4 88.7 87.3 91.0 90.8 88.8 88.6 89.3 SDv1.4 89.9 91.5 91.1 85.3 88.4 67.7 85.2 89.8 86.1 SDv1.5 91.6 91.2 91.1 85.8 91.0 87.9 88.8 90.3 89.7 ADM 86.4 85.7 85.9 87.0 87.1 86.8 86.9 85.2 86.4 Glide 89.3 87.0 86.9 89.8 91.4 90.8 90.4 87.5 89.1 Wukong 85.4 78.1 78.3 88.0 93.4 93.1 91.3 80.8 86.1 VQDM 86.9 86.7 86.7 88.3 89.0 88.6 88.5 86.1 87.6 BigGAN 89.8 89.6 89.5 86.3 90.5 91.5 90.3 89.5 89.6 Avg. 88.8 87.3 87.3 87.2 90.2 87.2 88.8 87.2 88.0
Table 2: Cross-generator evaluation of $k$-NN on the GenImage dataset for fake image detection. Accuracy of $k$-NN ($k=45$) classifier with support set size of 1,000 using distance correlation.
@@
CNNSpot [wang2020cnn] 58.2 70.3 70.2 57.0 57.1 67.7 56.7 56.6 61.7
Spec [zhang2019detecting] 56.7 72.4 72.3 57.9 65.4 70.3 61.7 64.3 65.1
F3-Net [qian2020thinking] 55.1 73.1 73.1 66.5 57.8 72.3 62.1 56.5 64.6
GramNet [liu2020global] 58.1 72.8 72.7 58.7 65.3 71.3 57.8 61.2 64.7
DIRE [wang2023dire] 65.0 73.7 73.7 61.9 69.1 74.3 63.4 56.7 67.2
LaRE2 [luo2024lare] 66.4 87.3 87.1 66.7 81.3 85.5 84.4 74.0 79.1
LATTE [vasilcoiu2025latte] 71.3 79.3 81.8 82.2 92.8 82.0 82.9 87.8 82.5
FRIDA-200 (Ours) 86.9 86.2 86.1 85.0 87.9 87.5 86.9 85.9 86.5
FRIDA (Ours) 88.8 87.3 87.3 87.2 90.2 87.2 88.8 87.2 88.0
Table 3: Comparison with state-of-the-art methods for fake image detection on the GenImage test set. Given a specified test set, we evaluate eight $k$-NN classifiers—each based on the support set from one generator—and compute their average accuracy. FRIDA-200 exploits a support set of 200 samples, reduced by $80\%$ w.r.t. FRIDA. Best in bold, second best underlined. † results taken by [luo2024lare].
As detailed in Table˜3, our approach demonstrates a significant leap in performance over current state-of-the-art methods on the GenImage test set. The experimental setup is designed to rigorously test generalization: eight models are trained on limited data (1,000 samples) from a single generator each, but evaluated on images from all generators. In this challenging scenario, our method is the one that better generalizes across these unseen data distributions, establishing a new state-of-the-art performance by a margin of nearly six percentage points.
Beyond its accuracy, the primary advantage of FRIDA in fake image detection lies in its remarkable data and training efficiency. The strong results are achieved without any conventional training process: we use a pre-trained SDM, and the $k$-NN framework merely requires storing the support set. The method also maintains high performance even with a drastically smaller support size. When the support set is reduced by $80\%$ (200 samples, referred to as FRIDA-200 in Table˜3), the average cross-generator accuracy on the test set drops by only $1.5\%$ (from $88.0\%$ to $86.5\%$), still outperforming the previous state-of-the-art (Table˜3). In a landscape where new generative models continually emerge, this property is crucial for practical deployment, as it avoids data-intensive retraining cycles.
@@ 5.3 Source Model Attribution
We compare MLP and $k$-NN classifiers on their ability to identify an image’s source generator, following the experimental procedure described in Section˜4.3.
@@
We adapt the distance-based approach used for fake image detection for the task of source model attribution. A portion of the results from the hyperparameters selection on the validation set obtained using the correlation distance is presented in Table˜4.These results prove that the $k$-NN classifier is inadequate for effective source model attribution. Despite optimization (using a support set of 9,000 and $k=9$), its peak accuracy reached only $57.7\%$, a performance level far too low for practical application.
@@
We train the three MLP classifiers as described in Section˜4.3; in Table˜5 we report the average results of the MLPs on the validation set across 10 training runs.
$k$ Support sizes 18 90 180 900 1800 3600 9000 1 25.2% 33.5% 36.4% 44.6% 48.1% 51.6% 55.8% 9 28.7% 31.3% 44.09% 48.7% 52.7% 57.7% 19 27.2% 42.5% 48.3% 52.1% 57.1% 35 39.5% 46.5% 50.9% 56.1% 45 37.8% 45.1% 50.2% 55.4%
Table 4: Effect of $k$ and support set size on source model attribution accuracy. Validation accuracies for the nine-class attribution task.
@@
Table 5: Source model attribution performance of MLPs. Average validation accuracy and standard deviation over 10 runs.
@@
While the performance of the three models on the source model attribution task are comparable, the MLP-640, with $84.87\%$ of average accuracy, slightly outperforms the others and allows for an increase in performance of about 27 percentage points over the $k$-NN classifier. The MLP-640, when tested on the GenImage test set, achieves an accuracy of $84.36\%$. This result is consistent with the validation performance, indicating a high degree of generalization. As shown in Figure˜3, the classifier generally recognizes source generators with high accuracy; however, it struggles to distinguish models that share the same architecture. Specifically, the model tends to confuse SDv1.4 with SDv1.5 and vice versa, while occasionally misattributing the SDM-based Wukong model to the same group. We attribute this phenomenon to their common architectural family. To validate this hypothesis, we conducted a family-based analysis by grouping data from these related sources, achieving a source model attribution accuracy of 96.67%.
@@
Notably, the huge performance increase of the MLP over the $k$-NN classifier on the validation set confirms that the features extracted by the diffusion U-Net contain generator-specific patterns. The MLP is capable of learning these distinctive characteristics, whereas the $k$-NN approach, which relies solely on feature distances, cannot recognize these subtle patterns. Although $k$-NN proves effective for the broader task of distinguishing real from synthetic content, it is insufficient for the more complex task of source model attribution. These results suggest that, while real and synthetic images can be detected through distances in the feature space, the discriminative power in source model attribution stems from more complex feature patterns that need an MLP to be captured. To investigate this hypothesis, we employ the SHAP algorithm to interpret the decisions of the MLP-640 classifier. We utilize the Gradient Explainer with a background dataset of 200 samples and 1,000 test samples. For each class, we identify the top 10 most informative features. In Figure˜4 we report the percentage of the top 10 features shared between each pair of generators.
Figure 3: Confusion Matrix for source model attribution. The MLP-640 is evaluated on the GenImage test set.
Figure 4: Shared informative features across generators. Percentage of overlap among the top 10 SHAP features.
@@
The two SDMs exhibit a $50\%$ overlap in their top 10 most important features. Furthermore, these shared features influence the model’s output identically (see LABEL:fig:shap_sdv14 and LABEL:fig:shap_sdv15 in the Supplementary Material), indicating that the images generated by these two models lead to a latent representation with common characteristics. As a result, the images generated by the two models are nearly indistinguishable to MLP. In contrast, the classifier successfully discriminates real and BigGAN images from the SDMs by exploiting divergent feature behavior. Although these models share $30\%$ of their top 10 features with the SDMs, the impact of these common features is markedly different (LABEL:fig:shap_real and LABEL:fig:shap_BigGAN in the Supplementary Material). Finally, the Wukong model presents an interesting paradox. Despite occasional misclassifications with the SDM-generated images, the SHAP analysis reveals zero feature overlap within the top 10 most influential features (LABEL:fig:shap_wukong in the Supplementary Material). This indicates that the classifier’s errors are not driven by the same high-impact features that define the SDM class. Instead, the confusion likely arises because a different set of features in Wukong images produces a combined effect that coincidentally mimics the characteristics of an SDM, leading the classifier to an incorrect conclusion.
5.4 Extended Experimental Analysis
Backbone Comparison.
In Table˜6 we report the results of the comparison among SDM, CLIP, and DINO as feature extractors following the setup described in Section˜4.4. Features from SDM consistently outperform features from the ViT encoders, validating the assumption that diffusion-based features are inherently more effective at isolating the discriminative cues necessary for distinguishing real from generated imagery.
Robustness to Image Perturbations.
In Figure˜5 we evaluate robustness under JPEG compression, Gaussian noise and Gaussian blur. While accuracy remains stable for JPEG factors up to 30 (a), it shows a steady decay under Gaussian noise (b). Notably, Gaussian blur (c) induces a sharp initial drop at $\sigma=1$ followed by a plateau, suggesting that the model successfully pivots to structural cues when local textures are suppressed.
Out-of-Distribution Robustness.
To evaluate the generalization capability of our method, we test FRIDA on several OOD datasets. Our approach achieves a solid average accuracy of 73.5%, demonstrating consistent detection across diverse generative architectures. Specifically, the model yields 69.4% on Flux 1.0 Dev, 68.1% on Flux 1.1 Pro, 72.5% on GPT Image 1, and 73.0% on SDv3.5. We observe a significant performance peak on Qwen (95.3%), while even on the most challenging generator in our benchmark, Ideogram 3.0, the detector maintains an accuracy of 62.7%.
Figure 5: Robustness analysis against image distortions. We report the detection accuracy for different image perturbations: (a) JPEG compression, (b) Gaussian noise, and (c) Gaussian blur.
Backbone Midj. SDv1.4 SDv1.5 ADM GLIDE BigGAN VQDM Wukong Avg
SD (Ours) 88.8 87.3 87.3 87.2 90.2 87.2 88.8 87.2 88.0
CLIP 63.6 64.6 64.5 57.9 66.3 60.2 55.9 64.1 61.5
DINO 58.0 62.5 62.3 57.5 66.2 63.0 61.9 63.7 61.2
Table 6: Backbone comparison. $k$-NN fake image detection replacing SDM with ViT-based backbones (CLIP and DINO).